Circular No. 6865

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J0421+560 AND CI CAMELOPARDALIS M. R. Garcia, P. Berlind, E. Barton, and J. E. McClintock, Harvard-

Smithsonian Center for Astrophysics; and P. J. Callanan and J. McCarthy, University College, Cork, report: “Photometric and spectroscopic observations of CI Cam, the possible counterpart of XTE J0421+56 (IAUC 6855, 6857), were obtained with the 1.2-m and 1.5-m telescopes at the Whipple Observatory during Apr. 3.08–3.17 UT. Conditions were photometric, with variable 2′′ seeing. We find V = 9.25 ± 0.1, B = 10.25 ± 0.1. This implies a brightening of > 2 mag in each band, relative to the levels measured by Bergner et al. (1995, A.Ap. Suppl. 112, 221). Relative photometry of three B and four V frames yields an upper limit to any short-term photometric variability of < 2 percent during our observations. Our spectra (range 400– 700 and 600–700 nm; resolution 0.3 and 0.075 nm), appear similar to those of Downes (1984, PASP 96, 807), but the level of Fe II and He I emission has strengthened in comparison to H Balmer. He II (468.53 nm) and He I (471.29 nm) are detected with equivalent widths of 1.5 ± 0.5 and 3.0 ± 1.0 nm, respectively, confirming the new He II emission reported by Wagner et al. (IAUC 6857). The forest of strong H, He I, and Fe II emission lines obscures the continuum to the extent that it is unclear if any photospheric absorption lines are present. None of the lines in our spectra shows any evidence for the type of double-peaked emission typically arising in x-ray-binary or cataclysmic-variable accretion disks. Our data are consistent with the association of CI Cam with XTE J0421+56, but imply that this might then be an unusual symbiotic-type x-ray binary, perhaps akin to GX 1+4 or 2A 1704+241 (= HD154791).”

GRB 980329 B. E. Schaefer, Yale University, writes: “R-band CCD exposures of

the error circle for GRB 980329 (IAUC 6854) were taken with the 3.5-m WIYN telescope on Apr. 1.17 and 3.10 UT. At the position of the radio tran-

sient (Taylor et al., http://gcn.gsfc.nasa.gov/gcn/gcn3/040.gcn3), no

source was detected, with limits of R > 24.2 and > 23.9, respectively.”

SUPERNOVA 1998S IN NGC 3877 Visual magnitude estimates: Mar. 28.85 UT, 12.4 (J. Shanklin, Cam-

bridge, England); 30.98, 12.5 (M. Suarez, Las Palmas, Canary Islands, Spain); Apr. 1.17, 12.5 (C. E. Spratt, Victoria, BC); 2.82, 12.7 (K. Hornoch, Lelekovice, Czech Republic); 3.97, 12.8 (M. Lehky, Veseli, Czech Rep.).

1998 April 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green